OpenStudy (anonymous):

Find the Domain of g(x) = √15-4 my answer = {x = all real numbers} right? because im not asked for the restriction...?

OpenStudy (m):

that's just a constant

OpenStudy (anonymous):

then? what else would i need to add? the restriction too?

OpenStudy (m):

g(x) = √15-4 is just a constant. there's no domain to input

OpenStudy (m):

is that what they gave you?

OpenStudy (anonymous):

it says find the domain, and then √15-4X ahh sorry i 4got the X

OpenStudy (m):

anything under radical cannot be a negative number

OpenStudy (m):

domain is X > 15/4

OpenStudy (m):

15/4 < X < infinity

OpenStudy (m):

(15/4, infinity)

OpenStudy (anonymous):

{X = all real numbers, 15/4 < X < infinity}

OpenStudy (m):

whoops i mean [15/4, infinity) since you can have 0 under radical

OpenStudy (m):

\[15/4 \le X < \infty\]

OpenStudy (anonymous):

i would just put that into brackets then right? like {X = all real numbers, 15/4 < X < ∞}

OpenStudy (m):

yeah but it's 15/4 ≤ X < ∞

OpenStudy (m):

greater than or equal to 15/4

OpenStudy (anonymous):

kk thanks